A car moving with a speed of 180km/hr was brought unformily to rest by the application of brakes in 20s. How far did the car travel after the brakes were applied?

Grade:12th pass

1 Answers

aditya shankar
3 years ago
apply formula              u=180km/hr=180*5/18=50m/s  t=20s  v=50-a*20 since final velocity is zero so, 0=50-20a ,a=2.5m/s
s=ut+.5gt^2
s=50*20-0.5*2.5*20*20
 =1000-500
=500m